# Charting the Landscape of African NLP: Mapping Progress and Shaping the Road Ahead

### v2

@@ Abstract
With over 2,000 languages and potentially millions of speakers, Africa represents one of the richest linguistic regions in the world. Yet, this diversity is scarcely reflected in state-of-the-art natural language processing (NLP) systems and large language models (LLMs), which predominantly support a narrow set of high-resource languages. This exclusion not only limits the reach and utility of modern NLP technologies but also risks widening the digital divide across linguistic communities. Nevertheless, NLP research on African languages is active and growing. In recent years, there has been a surge of interest in this area, driven by several factors—including the creation of multilingual language resources, the rise of community-led initiatives, and increased support through funding programs. In this survey, we analyze 734 research papers on NLP for African languages published over the past five years, offering a comprehensive overview of recent progress across core tasks. We identify key trends shaping the field and conclude by outlining promising directions to foster more inclusive and sustainable NLP research for African languages.
@@ 1 Introduction
    Identification of Open Issues: Through an extensive discussion of the current body of work, we highlight critical open issues—such as the imbalance in resources across African languages and the need for non-translated, native-language datasets—which can guide the strategic development of language technologies tailored to Africa [^1] We will release the collected dataset upon acceptance..
Figure 2: Inclusion flow of the systematic review.
@@ 2 Related Surveys
African languages are usually classified as low-resource languages. In their six-class classification scheme, Joshi et al. (2020) placed most African languages into the classes "left-behinds", scraping-bys", and "hopefuls" due to limited or nonexistent labeled data. Surveys exist for low-resource languages in general, covering common NLP approaches to boost performance in settings with little data (Hedderich et al., 2021; Haddow et al., 2022; Chen et al., 2023). Nigatu et al. (2024) argue however that the definition of "low-resource" is more complex than just a measurement of data availability, including also the availability of other resources and socio-political factors. This paper, therefore, focuses specifically on languages on the African continent which could have substantially different settings compared to low-resource languages in other parts of the world.
Several studies have focused on languages within specific countries, such as South Africa (Grover et al., 2010), Ethiopia (Tonja et al., 2023), Kenya (Amol et al., 2024), and Nigeria (Nwafor and Andy, 2022; Inuwa-Dutse, 2025). These surveys provide insights into the linguistic landscape, document available language resources, and explore computational methodologies tailored to the languages spoken in these countries. By offering a country-level perspective, they contribute to a broader understanding of language preservation and technological advancements within specific national contexts. Beyond country-specific studies, language-specific surveys have been conducted for individual African languages such as southern Sotho (Sibeko and Setaka, 2022) and Yoruba (Jimoh et al., 2025). However, a survey looking at the broader African continent is necessary to uncover both unique and shared challenges of its languages, and chart a cohesive future for building technologies for African languages.
@@ 3 Survey Methodology
To conduct a comprehensive systematic survey of the relevant literature, we employ a multi-source approach that integrates automated and expert-driven methods. First, we use the Semantic Scholar API, which retrieves papers by matching query terms against titles and abstracts in its database. The inclusion criteria are “Africa”, the names of the 54 African countries, and 53 African languages from NLLB-200 (NLLB Team et al., 2024), selected for their linguistic diversity and available NLP resources. These searches are carried out at the top conferences, journals, and workshops on NLP, speech, HCI, and AI conferences, journals, and workshops, as well as the preprint repository arXiv, allowing us to collect 1,454 relevant publications for a range of 5 years from 2019 to November 2024, and we provide the full list of venues in Appendix A. Next, we enhance this dataset by incorporating expert recommendations, ensuring the inclusion of influential and high-impact works that the Semantic Scholar API fails to retrieve, adding 230 other publications. By combining these sources, our approach ensures a well-rounded literature survey that balances algorithmic discovery with human expertise, leading to a more comprehensive and accurate understanding of the topic.
@@
The automatically retrieved articles contained irrelevant articles not focused on AI or more specifically, NLP research, especially from HCI venues and pre-prints. To address this, we filtered the papers by prompting GPT-4o to classify whether they were AI-related based on their title and abstract. Details for reproducibility are provided in Appendix B.
@@
We manually coded all the extracted papers with a codebook inspired by the *ACL tracks and iteratively refined during the coding process. We identify the languages covered, the NLP tasks addressed, the low-resource techniques employed, and other relevant themes. In addition, our annotation process included determining whether the paper introduced a dataset and, if so, whether it was released, cross-cultural, or translated from one language to another. We also assessed whether the paper proposed a model and, if so, whether the model was released. After the annotation process. We excluded papers on African American Vernacular English (AAVE) and those on spoken English that is not African-accent. Figure 2 shows an overview of the data filtering process.
Resulting Publication Set
In summary, we analyze a total of 734 papers. These include papers from top ACL conferences and workshops, as well as those recommended by domain experts.
@@ 4.1 How much do researchers publish?
Figure 5 illustrates the distribution of the papers over the five-year period under review. The number of publications grew significantly—from approximately 47 papers in 2019 to 101 in 2020—continuing to rise and peaking at more than 207 papers in 2023. A slight decline was observed in 2024, likely due to the cutoff for data collection in November. We hypothesize that this overall growth is driven by three key factors: (1) increased efforts to develop multilingual language resources, including embeddings and language models; (2) the emergence of community-led initiatives such as MasakhaneNLP [^2] https://www.masakhane.io/, and EthioNLP [^3] https://ethionlp.github.io/; and (3) the support of targeted funding programs such as Lacuna Fund [^4] https://lacunafund.org/ and AI4D [^5] https://www.ai4d.ai/, which have provided crucial resources for African language research. Furthermore, shared tasks have contributed to this momentum, with one paper in 2019, none in 2020, eight in 2021, 12 in 2022, 35 in 2023, and 21 in 2024. This upward trend highlights the growing and sustained interest in African NLP research.
@@ 4.2 What languages do researchers work on?
Given the annotated data, we evaluate all the African languages covered by the publications, using the African languages listed in Glottolog [^6] https://github.com/glottolog/glottolog-cldf/blob/master/cldf/languages.csv (Hammarström et al., 2024) as a reference. Our analysis shows that a total of 2,271 African languages and dialects—including dialects of Arabic, creoles, and sign languages from various African countries—have been covered by the papers. However, only about 20% of these languages appear in at least 10 papers, while over 70% were featured in less than 5 papers, highlighting the skewed distribution of language coverage. Figure 5 presents the top 20 languages by frequency, which closely align with the rankings of the most spoken languages in Africa according to LinguaMeta (Ritchie et al., 2024). In particular, 11 of the top 20 languages according to publication counts appear among the most widely spoken languages. However, Afrikaans (7.3M speakers) and other South African languages such as Tswana (6M), Pedi (5.3M), and Tsonga (2.5M) appeared among the top 20 languages covered in the papers—demonstrating the significant research efforts dedicated to South African languages despite their relatively lower positions in the LinguaMeta ranking. In terms of language families, Figure 5 shows that Atlantic–Congo languages dominate the papers, followed closely by Afro-Asiatic, with other families receiving much less coverage.
In addition to indigenous African languages, some papers, especially those with a focus on speech also included African accents of widely spoken non-indigenous African languages such as French, English, and Portuguese. While some papers clearly specified the country-specific accents they addressed (Aryal et al., 2023; Afonja et al., 2021; Muhammad et al., 2023; Hagemeijer et al., 2022a), others were more generic (Shan et al., 2023) in their approach. Furthermore, African languages are often code-mixed (Bandia, 1996), typically with English or other languages spoken within the same region. Our analysis shows that 5.5% of the reviewed articles focused on code mixing, with emphasis on the dialects of the North African Arabic dialects (Hamed et al., 2023, 2024), and South African languages (Wilkinson et al., 2020), but also several other cases (Diallo et al., 2021; Muhammad et al., 2022; Ilevbare et al., 2024).
@@
Overall, in terms of continent coverage, Figure 1 shows that every country has at least one language covered, with Mozambique and Nigeria leading at 347 and 320 papers respectively. Uganda, Tanzania, Rwanda, Burundi and Benin also have strong representation, each with over 280 papers. This shows notable skew in the geographic distribution of NLP research, with certain countries receiving higher attention.
@@ 4.3 What tasks do researchers work on?
The papers cover a skewed distribution of NLP tasks, with many addressing multiple tasks rather than just one. For instance, a paper that introduces a language model and evaluates it on a named entity recognition task contributes to both language modeling and named entity recognition. Figure 6 shows the distribution of these tasks, and we highlight the 10 major tasks by publications below:
Natural Language Understanding (NLU) is the most common task category. Within this category, multiple subcategories of sequence classification tasks exist, such as text or topic classification (Adelani et al., 2023; Ma et al., 2023), hateful or offensive or abusive speech detection (Thirion et al., 2020), emotion detection (Moudjari et al., 2020; Martin et al., 2022), natural language inference (Lin et al., 2022; Ahuja et al., 2023; Adelani et al., 2025), intent detection (Moghe et al., 2023; Skiredj et al., 2024), slot filling (FitzGerald et al., 2023; Mastel et al., 2023), semantic parsing (Ruder et al., 2023), and sentiment analysis (Diallo et al., 2021; Muhammad et al., 2022; Shode et al., 2023). Sequence classification tasks appear to be more common, likely due to the relative ease of creating datasets for them. This is especially true with techniques like annotation projection (Adelani et al., 2023; Ma et al., 2023), which facilitate multilingual data creation when parallel corpora are available. While some works focus on building datasets, others concentrate on developing methods to address these tasks, often through shared task contributions. Notably, around 45% of the NLU papers focus on sentiment analysis.
@@
Machine Translation (MT), the second most common task category, includes work on bitext mining and alignment (Heffernan et al., 2022; Feng et al., 2022)—especially from web sources—which has enabled the creation of widely used datasets like JW300 (Agić and Vulić, 2019), CCAligned (El-Kishky et al., 2020), and WikiMatrix (Schwenk et al., 2021), despite some concerns about content quality (Kreutzer et al., 2022). Multilingual benchmarks like NTREX-128 (Federmann et al., 2022), Flores (Goyal et al., 2022; NLLB Team et al., 2024), MAFAND-MT (Adelani et al., 2022a), and AfroLingu-MT (Elmadany et al., 2024) have also incorporated African languages. Furthermore, several efforts target specific language pairs (Sánchez-Martínez et al., 2020; Nguer et al., 2020; Adelani et al., 2021b; Azunre et al., 2021). Much of the work focuses on training and evaluating MT systems, including LLMs, for translation between African and high-resource languages (Adelani et al., 2022a; Elmadany et al., 2024). Furthermore, standard metrics like BLEU are inadequate for these morphologically rich languages, leading to the development of embedding-based metrics like AfriCOMET (Wang et al., 2024a, b), though its coverage remains limited.
@@
Speech processing: Aside from text-based research, there is a substantial amount of work focused on speech modality, including automatic phoneme or speech recognition (ASR) (Dossou and Emezue, 2021; Mohamud et al., 2021; Conneau et al., 2022; Sikasote et al., 2023a; Ogunremi et al., 2024), text-to-speech (TTS) (Ogayo et al., 2022; Meyer et al., 2022; Ma et al., 2024), speaker recognition (Villalba et al., 2019), tone recognition (Obiang et al., 2024), emotion recognition, speech translation (Zanon Boito et al., 2022; Sikasote et al., 2023a; Ahia et al., 2024), and acoustic modeling (Vyas et al., 2020; Tachbelie et al., 2020). Recently, there has been growing interest in speech representation learning, particularly with transformer-based models. A few models include African languages, both African-centric (Kimanuka et al., 2024; Caubrière and Gauthier, 2024; Alabi et al., 2024b) and massively multilingual models (Conneau et al., 2020a; Zanon Boito et al., 2024; Chen et al., 2024c). These advances are driven by efforts in the massive curation of speech resources. However, despite the broad coverage of these models, the most commonly used evaluation benchmark for ASR, FLEURS (Conneau et al., 2022), covers only 21 African languages, creating an imbalance in evaluation.
@@
Language models: During the survey period, there was a notable increase in the development of large transformer-based language models. Many studies introduced such models for African languages, including monolingual models for individual languages (Nzeyimana and Niyongabo Rubungo, 2022; Martin et al., 2022), and massively multilingual models that include African languages alongside others (Conneau and Lample, 2019; Imani et al., 2023). These models were developed by training them from scratch (Jude Ogundepo et al., 2022; Ogueji et al., 2021) or extending existing language models (Alabi et al., 2022; Adelani et al., 2024; Meyer et al., 2024). These advances, like those in speech, rely on large-scale language data from the web. In this line are also works on representation learning, which includes learning embedding representation for text (Yuan et al., 2020; Liu et al., 2023) or speech (Jacobs, 2024), and analyzing them (Alabi et al., 2020).
@@
Information extraction (IE): is another common category of NLP tasks and includes work on named entity recognition (NER) (Adelani et al., 2021a, 2022b; Rahimi et al., 2019; Oyewusi et al., 2021; Mbuvha et al., 2023), entity linking and typing (Lin et al., 2019; Zhu et al., 2019), event extraction (Jenkins et al., 2023), relation classification (Lent et al., 2024), and keyword spotting and localization (Yusuf and Saraçlar, 2019; Olaleye et al., 2022; Nortje et al., 2024). Among these, NER is particularly common due to the availability of resources such as MasakhaNER (Adelani et al., 2021a, 2022b) and Wikiann NER (Pan et al., 2017).
@@
Question Answering (QA)—which involves extracting, retrieving, or generating answers to input queries—remains an emerging research area for African languages. Recent efforts have focused on creating benchmarks from sources such as Wikipedia, for both QA and machine reading comprehension across multiple languages. For example, multilingual datasets like TyDi QA (Clark et al., 2020), which includes only Swahili from Africa, AfriQA (Ogundepo et al., 2023), and NaijaRC (Aremu et al., 2024). There have also been language-specific initiatives (Wanjawa et al., 2023; Taffa et al., 2024). Another relevant work is Belebele (Bandarkar et al., 2024), a machine reading comprehension dataset translated into over 200 languages, including 50+ African languages.
Reasoning: Closely related to QA is the growing focus on evaluating the reasoning abilities of large language models. Recent work explores various types of reasoning—linguistic (Bean et al., 2024), cultural knowledge (Myung et al., 2024a), commonsense (Ponti et al., 2020), mathematical (Adelani et al., 2025), ethical (Agarwal et al., 2024), and moral (Khandelwal et al., 2024). Notably, many of the datasets used for such evaluations in African languages are translated from English, which may limit cultural relevance and linguistic nuance. Evaluation results show that LLMs generally perform better on these tasks in high-resource languages compared to low-resource ones, including African languages (Adelani et al., 2025).
@@
Syntax, Tagging, and Parsing: A total of 59 papers have explored various syntax-related tasks, including dependency parsing (Seddah et al., 2020; Dione, 2021; Steimel and Kübler, 2023; Ralethe, 2020), syntactic parsing (Momoh, 2024), part-of-speech (POS) tagging (Dione et al., 2023; Faisal et al., 2024), and computational grammar (Hellan, 2020). These efforts span both resource creation and evaluation; however, they cover only a limited number of African languages.
Others: Other tasks categorization with less publications including semantics, multimodality (text + image or video), text generation (excluding machine translation) are described in Appendix C.

### v3

@@ Abstract
With over 2,000 languages and potentially millions of speakers, Africa represents one of the richest linguistic regions in the world. Yet, this diversity is scarcely reflected in state-of-the-art natural language processing (NLP) systems and large language models (LLMs), which predominantly support a narrow set of high-resource languages. This exclusion not only limits the reach and utility of modern NLP technologies but also risks widening the digital divide across linguistic communities. Nevertheless, NLP research on African languages is active and growing. In recent years, there has been a surge of interest in this area, driven by several factors—including the creation of multilingual language resources, the rise of community-led initiatives, and increased support through funding programs. In this survey, we analyze 884 research papers on NLP for African languages published over the past five years, offering a comprehensive overview of recent progress across core tasks. We identify key trends shaping the field and conclude by outlining promising directions to foster more inclusive and sustainable NLP research for African languages. [^1] We release our data and code publicly at https://github.com/uds-lsv/africanlp-survey
@@ 1 Introduction
    Identification of Open Issues: Through an extensive discussion of the current body of work, we highlight critical open issues—such as the imbalance in resources across African languages and the need for non-translated, native-language datasets—which can guide the strategic development of language technologies tailored to Africa.
@@ 2 Related Surveys
African languages are usually classified as low-resource languages. In their six-class classification scheme, Joshi et al. (2020) placed most African languages into the classes "left-behinds", scraping-bys", and "hopefuls" due to limited or non-existent labeled data. Surveys exist for low-resource languages in general, covering common NLP approaches to boost performance in settings with little data (Hedderich et al., 2021; Haddow et al., 2022; Chen et al., 2023). However, as Nigatu et al. (2024) point out, the definition of "low-resource" is more complex than just a measurement of data availability, including also the availability of other resources and socio-political factors. This paper therefore focuses specifically on languages on the African continent, which could have substantially different settings compared to low-resource languages in other parts of the world.
Several studies have focused on languages within specific countries, such as South Africa (Grover et al., 2010), Ethiopia (Tonja et al., 2023), Kenya (Amol et al., 2024), Nigeria (Nwafor and Andy, 2022; Inuwa-Dutse, 2025), and Ghana (Azunre et al., 2021a; Issaka et al., 2024). These surveys provide insights into the linguistic landscape, document available language resources, and explore computational methodologies tailored to the languages spoken in these countries. By offering a country-level perspective, they contribute to a broader understanding of language preservation and technological advancements within specific national contexts. Beyond country-specific studies, language-specific surveys have been conducted for individual African languages such as Southern Sotho (Sibeko and Setaka, 2022) and Yoruba (Jimoh et al., 2025). However, a survey looking at the broader African continent is necessary to uncover both unique and shared challenges of its languages, and chart a cohesive future for building technologies for African languages.
Similar to our effort, the survey by Mussandi and Wichert (2024) provides a tool-centric overview of NLP resources available for African languages. Keet (2022) instead reviews how bootstrapping has been applied in practice to the Niger-Congo (‘Bantu’) subset of African languages, showing that rule- and grammar-based approaches transfer more successfully across distant languages, whereas data-driven methods work reliably only for closely related languages. In contrast, Ikae and Kurpicz-Briki (2024) concentrates on bias detection in machine translation for African and European languages. Our work differs by offering a broader landscape analysis of African NLP research, highlighting trends in language coverage, tasks, and research output.
@@ 3 Survey Methodology
To conduct a comprehensive systematic survey of the relevant literature, we employ a multi-source approach that integrates automated and expert-driven methods. First, we use the Semantic Scholar API, which retrieves papers by matching query terms against titles and abstracts in its database. The inclusion criteria are “Africa”, the names of the 54 African countries, and 2,290 African languages from the Glottolog database. [^2] Glottolog documents over 3,000 African languages and dialects, but we used only those with an official ISO 639-3 code. These searches are conducted across leading conferences, journals, and workshops in NLP, speech, HCI, and AI, as well as the preprint repository arXiv. The full list of venues is provided in Appendix˜A. We then augment this dataset with additional relevant papers to ensure the inclusion of influential and high-impact works that the Semantic Scholar API does not capture, adding 209 publications. In total, this process yields 27,593 relevant publications spanning five years, from 2019 to 2024. By combining these sources, our approach ensures a well-rounded literature survey that balances algorithmic discovery with human expertise, leading to a more comprehensive and accurate understanding of the topic.
@@
The automatically retrieved articles contained irrelevant articles not focused on AI or more specifically, NLP research, especially from HCI venues and pre-prints. To address this, we filtered the papers by prompting GPT-4o to classify whether they were NLP-related based on their title and abstract. Details for reproducibility are provided in Appendix˜B.
@@
We manually coded all the extracted papers with a codebook inspired by the ACL 2025 tracks and iteratively refined during the coding process. We identify the languages covered, the NLP tasks addressed, the low-resource techniques employed, and other relevant themes. In addition, our annotation process included determining whether the paper introduced a dataset and, if so, whether it was released and whether it was cross-cultural or translated from one language to another. We also assessed whether the paper proposed a model and, if so, whether the model was released. We excluded papers on African American Vernacular English (AAVE) and those on spoken English that is not African-accent. We include African-accented variants of European languages because they are part of the linguistic reality of the continent, and handling them well is crucial for making NLP inclusive for Africans. Figure˜2 shows an overview of the data filtering process. The resulting publication set contains a total of 884 papers.
@@ 4.1 How much do researchers publish?
Figure˜5 illustrates the distribution of the papers over the five-year period under review. The number of publications grew significantly—from approximately 57 papers in 2019 to 117 in 2020—continuing to rise and peaking at 232 papers in 2024. We hypothesize that this overall growth is driven by three key factors: (1) increased efforts to develop multilingual language resources, including embeddings and language models; (2) the emergence of community-led initiatives such as MasakhaneNLP, [^3] https://www.masakhane.io/ EthioNLP, [^4] https://ethionlp.github.io/ HausaNLP, [^5] https://hausanlp.github.io/ and, GhanaNLP; [^6] https://ghananlp.org/ and (3) the support of targeted funding programs such as Lacuna Fund [^7] https://lacunafund.org/ and AI4D [^8] https://www.ai4d.ai/, which have provided crucial resources for African language research. Furthermore, shared tasks have contributed to this momentum, with one paper each in 2019 and 2020, 11 in 2021, 15 in 2022, 41 in 2023, and 24 in 2024. This upward trend highlights the growing and sustained interest in African NLP research.
@@ 4.2 What languages do researchers work on?
Given the annotated data, we evaluate all the African languages covered by the publications, using the African languages listed in Glottolog [^9] https://github.com/glottolog/glottolog-cldf/blob/master/cldf/languages.csv (Hammarström et al., 2024) as a reference. Our analysis shows that a total of 2,275 African languages and dialects—including dialects of Arabic, creoles, and sign languages from various African countries—have been covered by the papers. However, only about 25% of these languages appear in at least 10 papers, while over 67% were featured in less than 5 papers, highlighting the skewed distribution of language coverage. Figure˜5 presents the top 20 languages by frequency, which closely align with the rankings of the most spoken languages in Africa according to LinguaMeta (Ritchie et al., 2024). In particular, 11 of the top 20 languages according to publication counts appear among the most widely spoken languages. However, Afrikaans (7.3M speakers) and other South African languages such as Tswana (6M), and Tsonga (2.5M) appeared among the top 20 languages covered in the papers—demonstrating the significant research efforts dedicated to South African languages despite their relatively lower positions in the LinguaMeta ranking. In terms of language families, Figure˜5 shows that Atlantic–Congo languages dominate the papers, followed closely by Afro-Asiatic, with other families receiving much less coverage.
In addition to indigenous African languages, some papers, particularly those focusing on speech also included African accents of widely spoken non-indigenous languages such as French, English, and Portuguese. While some studies clearly specified the country-specific accents they addressed (Aryal et al., 2023; Afonja et al., 2021; Olatunji et al., 2023b; Hagemeijer et al., 2022a), others were more generic (Shan et al., 2023). There are also research efforts on texts in non-indigenous languages tailored to African contexts, such as AfriSenti (Muhammad et al., 2023) and AfriMed-QA (Nimo et al., 2025).
@@
Furthermore, African languages are often code-mixed (Bandia, 1996), typically with English or other languages spoken within the same region. Our analysis shows that 5.1% of the reviewed articles focused on code mixing, with emphasis on the dialects of the North African Arabic dialects (Hamed et al., 2023, 2024), and South African languages (Wilkinson et al., 2020), but also several other cases (Diallo et al., 2021; Muhammad et al., 2022; Ilevbare et al., 2024).
Overall, in terms of continent coverage, Figure˜1 shows that every country has at least one language covered, with Mozambique and Uganda leading at 445 and 383 papers respectively. Tanzania, Nigeria, Rwanda, Burundi, Kenya and Somalia also have strong representation, each with over 350 papers. This shows a notable skew in the geographic distribution of NLP research, with certain countries receiving more attention.
@@ 4.3 What tasks do researchers work on?
The papers cover a skewed distribution of NLP tasks, with many addressing multiple tasks rather than just one. For instance, a paper that introduces a language model and evaluates it on a named entity recognition task contributes to both language modeling and named entity recognition. Figure˜6 shows the distribution of these tasks, and we highlight the 10 major tasks by publications below:
Natural Language Understanding (NLU) is the most common task category. Within this category, multiple subcategories of sequence classification tasks exist, such as text or topic classification (Adelani et al., 2023; Ma et al., 2023), hateful or offensive or abusive speech detection (Thirion et al., 2020), emotion detection (Moudjari et al., 2020; Martin et al., 2022), natural language inference (Lin et al., 2022; Ahuja et al., 2023; Adelani et al., 2025), intent detection (Moghe et al., 2023; Skiredj et al., 2024), slot filling (FitzGerald et al., 2023; Mastel et al., 2023), semantic parsing (Ruder et al., 2023), and sentiment analysis (Diallo et al., 2021; Muhammad et al., 2022; Shode et al., 2023). Notably, around 40% of the NLU papers focus on sentiment analysis. Sequence classification tasks appear to be more common, likely due to the relative ease of creating datasets for them. This is especially true with techniques like annotation projection (Adelani et al., 2023; Ma et al., 2023), which facilitate multilingual data creation when parallel corpora are available. While some works focus on building datasets, others concentrate on developing methods to address these tasks, often through shared task contributions.
@@
Machine Translation (MT), the second most common task category, includes work on bitext mining and alignment (Heffernan et al., 2022; Feng et al., 2022)—especially from web sources—which has enabled the creation of widely used datasets like JW300 (Agić and Vulić, 2019), CCAligned (El-Kishky et al., 2020), and WikiMatrix (Schwenk et al., 2021), despite some concerns about content quality (Kreutzer et al., 2022). Multilingual benchmarks like NTREX-128 (Federmann et al., 2022), Flores (Goyal et al., 2022; NLLB Team et al., 2024), MAFAND-MT (Adelani et al., 2022a), and AfroLingu-MT (Elmadany et al., 2024) have also incorporated African languages. Furthermore, several efforts target specific language pairs (Sánchez-Martínez et al., 2020; Nguer et al., 2020; Adelani et al., 2021b; Azunre et al., 2021b). Much of the work focuses on training and evaluating MT systems, including LLMs, for translation between African and high-resource languages (Adelani et al., 2022a; Elmadany et al., 2024). Furthermore, standard MT metrics like BLEU are inadequate for these morphologically rich languages, leading to the development of embedding-based metrics like AfriCOMET (Wang et al., 2024a, b), though its coverage remains limited.
@@
Speech Processing: Aside from text-based research, there is a substantial amount of work focused on speech modality, including automatic phoneme or speech recognition (ASR) (Dossou and Emezue, 2021; Mohamud et al., 2021; Conneau et al., 2022; Sikasote et al., 2023a; Ogunremi et al., 2024), text-to-speech (TTS) (Ogayo et al., 2022; Meyer et al., 2022; Ma et al., 2024; Lux et al., 2024), speaker recognition (Villalba et al., 2019), tone recognition (Obiang et al., 2024), emotion recognition, speech translation (Zanon Boito et al., 2022; Sikasote et al., 2023a; Ahia et al., 2024), and acoustic modeling (Vyas et al., 2020; Tachbelie et al., 2020). Recently, there has been growing interest in speech representation learning, particularly with transformer-based models. A few models include African languages, both African-centric (Kimanuka et al., 2024; Caubrière and Gauthier, 2024; Alabi et al., 2025) and massively multilingual models (Conneau et al., 2020a; Zanon Boito et al., 2024; Chen et al., 2024c). These advances are driven by efforts in the massive curation of speech resources. However, despite the broad coverage of these models, the most commonly used evaluation benchmark for ASR, FLEURS (Conneau et al., 2022), covers only 21 African languages, creating an imbalance in evaluation.
@@
Language Models (LMs): During the survey period, there was a notable increase in the development of large transformer-based language models. Many studies introduced such models for African languages, including monolingual models for individual languages (Nzeyimana and Niyongabo Rubungo, 2022; Martin et al., 2022), and massively multilingual models that include African languages alongside others (Conneau and Lample, 2019; Imani et al., 2023). These models were developed by training them from scratch (Jude Ogundepo et al., 2022; Ogueji et al., 2021) or extending existing language models (Alabi et al., 2022; Adelani et al., 2024; Meyer et al., 2024). These advances, like those in speech, rely on large-scale language data from the web. In this line are also works on representation learning, which includes learning embedding representation for text (Yuan et al., 2020; Liu et al., 2023) or speech (Jacobs, 2024), and analyzing them (Alabi et al., 2020). In Section˜E.2, we provide an overview of several models for African languages, including LMs.
@@
Information Extraction (IE) is another common category of NLP tasks and includes work on named entity recognition (NER) (Adelani et al., 2021a, 2022b; Rahimi et al., 2019; Oyewusi et al., 2021; Mbuvha et al., 2023), entity linking and typing (Lin et al., 2019; Zhu et al., 2019), event extraction (Jenkins et al., 2023), relation classification (Lent et al., 2024), and keyword spotting and localization (Yusuf and Saraçlar, 2019; Olaleye et al., 2022; Nortje et al., 2024). Among these, NER is particularly common due to the availability of resources such as MasakhaNER (Adelani et al., 2021a, 2022b) and Wikiann NER (Pan et al., 2017).
@@
Question Answering (QA)—which involves extracting, retrieving, or generating answers to input queries—remains an emerging research area for African languages. Recent efforts have focused on creating benchmarks from sources such as Wikipedia, for both QA and machine reading comprehension across multiple languages. For example, multilingual datasets like TyDi QA (Clark et al., 2020), which includes only Swahili from Africa, AfriQA (Ogundepo et al., 2023), NaijaRC (Aremu et al., 2024), and AfriMed-QA (Nimo et al., 2025). There have also been language-specific initiatives (Wanjawa et al., 2023; Taffa et al., 2024; Costa-jussà et al., 2025). Another relevant work is Belebele (Bandarkar et al., 2024), a machine reading comprehension dataset translated into over 200 languages, including 50+ African languages.
Reasoning: Closely related to QA is the growing focus on evaluating the reasoning abilities of large language models. Recent work explores various types of reasoning—linguistic (Bean et al., 2024), cultural knowledge (Myung et al., 2024a), commonsense (Ponti et al., 2020), mathematical (Adelani et al., 2025), comparative (Agrawal et al., 2024), ethical (Agarwal et al., 2024), and moral (Khandelwal et al., 2024). Notably, many of the datasets used for such evaluations in African languages are translated from English, which may limit cultural relevance and linguistic nuance. Evaluation results show that LLMs generally perform better on these tasks in high-resource languages compared to low-resource ones, including African languages (Adelani et al., 2025).
@@
Syntax, Tagging, and Parsing: A total of 68 papers have explored various syntax-related tasks, including dependency parsing (Seddah et al., 2020; Dione, 2021; Steimel and Kübler, 2023; Ralethe, 2020), syntactic parsing (Momoh, 2024), part-of-speech (POS) tagging (Dione et al., 2023; Faisal et al., 2024), and computational grammar (Hellan, 2020). These efforts span both resource creation and evaluation; however, they cover only a limited number of African languages.
Others: Other tasks categorization with less publications including semantics, multimodality (text + image or video), text generation (excluding machine translation) are described in Appendix˜C.